# DOEPatch: Dynamically Optimized Ensemble Model for Adversarial Patches Generation

### v1

@@ II-A Ensemble-based Attacks
Early adversarial attack research primarily focused on digital world attacks against image classification models[31, 32, 33]. By adding imperceptible perturbations to the original images, classification models could be deceived and make incorrect predictions. The resulting images with added perturbation were known as adversarial examples. With the widespread application of adversarial examples, they could be transferred between different architectures, severely affecting the applicability of neural networks. Some efforts have been made to enhance the transferability of adversarial examples by training substitute models[34], but it was challenging to apply this approach to large models and large datasets. Liu et al.[35] pioneered the research on the transferability of large-scale models and datasets by introducing an approach based on the ensemble of multiple target models to generate transferable adversarial examples. Based on the ensembled model attack, by training an adversarial sample that can deceive multiple attacked models, the success rate of attacks against unknown black-box models will also increase. Zheng et al.[36] introduced the Meta Gradient Adversarial Attack (MGAA) framework, which can be seamlessly ensembled with existing gradient-based attack methods to improve the transferability of adversarial examples across different models. Despite the extensive research on adversarial attacks in the field of image classification, there has been a lack of attention on ensemble-based attacks for dense prediction models such as object detection[26, 27]. Meanwhile, most of the existing ensemble model works[37, 38, 39] simply fuse the outputs of all models, neglecting the differences in model structures, which leads to adversarial optimization mostly heading in a single direction, making it difficult for adversarial examples to converge into a form that is effective for all target models. Recently, Cai et al.[17] proposed an attack algorithm based on transfer and query combination, demonstrating that adjusting the weights of the ensemble according to the victim model can further improve the attack performance. The universal adversarial perturbations they obtained can successfully attack various object detection models, such as YOLOv3 and Faster R-CNN, but can only be applied to digital images. Our research focused on the application of ensemble models in adversarial patches, adjusting weight coefficients based on energy loss feedback from different victim models to generate effective patches for all models.
@@ III-A Problem Definitions
In contrast to traditional methods of generating adversarial samples, adversarial patches are like stickers that are placed over specific local regions of an image. In order to investigate the generation mechanism of adversarial patches, we introduce the concept of the energy model and redefine the process of generating adversarial patches for object detection models from an energy perspective. Generally, in energy-based model, the training set ${S}$ is comprised of tuples ${\{({X^{i}},{Y^{i}});i=1...,N\}}$, which ${{X^{i}}}$ represents the input training example and ${{Y^{i}}}$ denotes its corresponding ground-truth label. Each learning approach can be described as the process of finding parameters ${G\in\left\{{\cal G}\right\}}$ that minimize the energy loss function ${\min L(G,S)}$. However, our adversarial training differs from the training processes of object detection models or energy-based models. The size and distribution of pixel values in the adversarial patch serve as the parameters ${G}$ that require updating. The energy loss function for constructing adversarial patches on the energy-based model is defined by the following equation,
@@
${T(\cdot)}$ represents the transformation function responsible for adding the adversarial patch to the input image, ${p\in{\mathbb{R}^{H\times W\times 3}}}$ denotes the adversarial patch, and ${R(p)={\left\|p\right\|_{\mathbb{P}}}{\text{ = (}}\sum\limits_{i,j}{{{\left|{{p_{(i,j)}}}\right|}^{\mathbb{P}}}}{{\text{)}}^{1/\mathbb{P}}}}$ is the regularization term that restricts the size of each pixel ${{p_{(i,j)}}}$ within the adversarial patch by imposing a ${{l_{\mathbb{P}}}}$-norm constraint. During the adversarial patch training process, we optimize the confidence scores of candidate bounding boxes belonging to the “person” category, which are used as the objective for the optimization algorithm (i.e., the output of ${f(Y,T(p,X))}$). The energy loss function incorporates the confidence scores of all candidate boxes as part of the optimization process. The adversarial patch serves as the parameter to be updated iteratively until the energy of the target category is minimized through the optimization process.
@@ III-B Generating Adversarial Patch
Transformation Function: To facilitate the generalization of the adversarial patch from the digital to the physical world, we leverage the Expectation over Transformation(EoT)[41] framework and introduce a set of physical transformations ${T_{EoT}}$ when applying the patch to an object’s surface. These transformations incorporate variations in contrast, brightness, angles, and other physical environmental factors, thereby ensuring the robustness and effectiveness of the adversarial patch across diverse physical settings. However, in real-world physical environments, the scale of the patch tends to decrease as distance increases and the resulting imagery becomes more blurred. Building on the EoT framework, we introduce a novel distance transformation function. We employ a Gaussian kernel to blur the image, ensuring that the smaller the target object becomes, the larger the convolution kernel becomes, resulting in a more blurred adversarial patch image. Our goal is to develop an adversarial patch pattern that can be printed on a T-shirt to enable a person to evade object detection models. To achieve this goal, we introduce the Thin Plate Splines(TPS)[42] approach to simulate non-rigid deformations such as folds on the T-shirt’s surface. Using the aforementioned transformation methods, we devise a transformation function ${{T_{p}}(\cdot)}$ to apply the adversarial patch to the target image,
$\begin{gathered}{X_{adv}}={T_{p}}(p,X)=\hfill\\ {t_{EoT}}((1-{M_{p}})\cdot X+{t_{d}}({M_{p}}\cdot{t_{TPS}}(p)))\hfill\\ \end{gathered}$ (3)
${{M_{p}}\in{[0,1]^{d}}}$ represents the mask region of the adversarial patch on the target image, which is determined based on the bounding box obtained from the label. The ${{T_{p}}(\cdot)}$ function encompasses three main types of transformations, of which the first one involves applying TPS transformation(${{t_{TPS}}\in{T_{TPS}}}$) to the generated adversarial patch. This serves to introduce distortions, thereby simulating non-rigid deformations on clothing surfaces. Subsequently, based on the size of the bounding box in the input image, the distance is estimated and a distance transformation function(${{t_{d}}\in{T_{d}}}$) is applied to the generated adversarial patch. This enables the patch image to be blurred to varying degrees. In the last step, a regular physical transformation(${{t_{EoT}}\in{T_{EoT}}}$) is applied to the adversarial patch to replicate typical variations that occur in physical environments. The complete process is illustrated in Fig.3,
@@
Adversarial Loss: For contemporary object detection models, it is crucial to design an appropriate energy function that can effectively deactivate the detection of the person category by minimizing the energy. Given images ${X}$ and the ground-truth labels ${Y}$, contemporary object detectors take this information as input, and utilize detection models to generate bounding boxes’ information and corresponding confidence scores for the target category. The candidates with scores exceeding a predefined threshold are selected as the predicted results. Consequently, we incorporate the expected confidence scores of the output as a part of the energy function for adversarial patch optimization, aiming to reduce the target’s confidence scores by optimizing the adversarial patch. Specifically, we begin by randomly initializing the adversarial patch and utilizing the transformation function ${{T_{p}}(\cdot)}$ to apply the patch to the original image, which serves as the input for the object detection models. Previous research incorporated the maximum confidence score of the target category, which was output by the detection network, as a constituent element in the loss function. Unlike previous work[18, 43], we consider all ${N}$ bounding boxes output by the detection model as optimization targets during each iteration, taking into account changes of all confidence scores comprehensively. To reduce the computational cost, we select candidate boxes whose confidence scores for the category “person” are higher than the set threshold value ${\mu}$ as attack objects. Subsequently, we continually optimize and update the patch pixels through gradient descent. In the case of one-stage object detection models, such as YOLOv2, YOLOv3, the detection network directly outputs all bounding boxes and associated category information. We integrate confidence scores of candidate boxes relevant to the “person” category as a constituent of the energy function. For two-stage object detection models like Faster R-CNN, person-related bounding boxes are extracted through Region Proposal Network(RPN), incorporating both positive and negative sample information. Positive candidate boxes with confidence scores surpassing a threshold are selectively chosen as attack targets. Hence, we compute the confidence score of all bounding boxes in the adversarial sample that exceeds the threshold ${\mu}$ of the target detection network. This score is merged into the energy function used for training the adversarial patch. Through iterative optimization, the patch is fine-tuned to minimize the total energy loss, as defined in Eq.4. The generative algorithm, designed based on the energy function, optimizes the adversarial patch during training, thereby reducing energy and causing a malfunction in the target detector’s person detection capability. By training a highly complex adversarial patch, the energy of the remaining categories is elevated, leading to a reduction in the energy associated with the person category. Sec.IV-D of the experiment provides a more detailed explanation regarding the impact of the adversarial patch on the energy levels of various categories.
@@ III-C Dynamically Optimized Ensemble Model
To achieve person invisibility across multiple object detection models, the generated adversarial patch effectively impacts the predictive information of all detection algorithms, resulting in a significant reduction of the energy associated with the person category in the output. For ${K}$ object detection models, ${f(\cdot)}$ denotes its outputs and ${\mathbf{w}=[{{\text{w}}_{1}},{{\text{w}}_{2}},...,{{\text{w}}_{K}}]}$ denotes the weights used to adjust different models, subject to constraint ${\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}}=1}$. The ensemble model for the adversarial patch can be expressed as follows,
$\mathop{\min}\limits_{p}\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}\cdot{f_{i}}(p)}$ (5)
However, it is challenging to achieve an ideal balance point for the weights in object detection models with vastly different structures such as YOLO and Faster R-CNN, and overfitting is therefore prone to occur during the training process. Sec.IV-E of this paper provides a detailed analysis of this phenomenon. To address the aforementioned challenges, we propose a dynamically optimized ensemble model for multiple object detection models based on the Min-Max adversarial training approach. During the training process, the proportions of each detection model’s contribution to the overall energy output are dynamically adjusted through parameter tuning to achieve the universality of the adversarial patch, as calculated in Eq.6. The training procedure involves an outer minimization phase, with fixed parameters ${\mathbf{w}}$, where the adversarial patch is optimized through backpropagation to minimize the energy output of the ensemble model. And an inner maximization phase is performed with a fixed adversarial patch, using projected gradient descent to balance the weights ${\mathbf{w}}$ of different models and achieve the maximum energy output of the ensemble model. However, excessive parameter fluctuations can lead to oscillations during the training process, making it difficult to converge to an optimal adversarial patch. On the other hand, if the parameter fluctuations are too small, the loss may stop decreasing after reaching a certain point, failing to reach the desired level. In order to stabilize the process of the weights ${\mathbf{w}}$, the dynamic rate adjustment parameter $\nu$ is increased to control the speed of ${\mathbf{w}}$ change. The value of the parameter $\nu$ was verified in Sec.IV-A.
@@ IV-A Implementation Details
To conduct experiments on digital image attacks, we utilized the Inria Person dataset[46] consisting of 614 person images for training and 288 images for testing purposes. To evaluate the attack effectiveness and transferability of adversarial patches, we first measure their impact by computing the Average Precision (AP) values of the images with and without the patches. The larger the drop in AP, the stronger the attack effectiveness. The adversarial patches designed in this study were mainly intended to deceive mainstream one-stage object detection models, namely YOLOv2, YOLOv3, and YOLOv4. As for more complex two-stage models, these patches were found to result in failure to predict when applied to Faster R-CNN detection models with backbones based on VGG16 and ResNet50 architectures. All detection models were trained on the COCO dataset, with the output confidence threshold set to 0.5 and the Non-Maximum Suppression (NMS) threshold to 0.4 during the inference stage. We consider three different sets of object detection models to design ensemble models for the adversarial patches:
@@
The entire algorithm is designed in PyTorch and trained on an NVIDIA RTX 3090 24GB GPU cluster. The adversarial patches are set to a size of $3\times 300\times 300$, optimized using the Adam[47] optimizer, and the learning rate for external minimization optimization is set to 0.03. The weight coefficients were determined based on experience, $\alpha=0.01$, $\beta=0.165$, $\mu=0.4$ and the regularization coefficient was set $\gamma=0.1$. The internal dynamic rate adjustment parameter $\nu$ was determined based on experimental results on DOEPatch(YF), as shown in Fig.4. Validation showed that with $\nu=0.78$, the adversarial patch achieved good attack effects on both YOLOv2 and YOLOv3.

### v2

@@ II-A Ensemble-based Attacks
Early adversarial attack research primarily focused on digital world attacks against image classification models[7, 31, 32]. By adding imperceptible perturbations to the original images, classification models could be deceived and make incorrect predictions. The resulting images with added perturbation were known as adversarial examples. With the widespread application of adversarial examples, they could be transferred between different architectures, severely affecting the applicability of neural networks. Some efforts have been made to enhance the transferability of adversarial examples by training substitute models, but it was challenging to apply this approach to large models and large datasets. Liu et al.[33] pioneered the research on the transferability of large-scale models and datasets by introducing an approach based on the ensemble of multiple target models to generate transferable adversarial examples. Based on the ensembled model attack, by training an adversarial sample that can deceive multiple attacked models, the success rate of attacks against unknown black-box models will also increase. Zheng et al.[34] introduced the Meta Gradient Adversarial Attack (MGAA) framework, which can be seamlessly ensembled with existing gradient-based attack methods to improve the transferability of adversarial examples across different models. Despite the extensive research on adversarial attacks in the field of image classification, there has been a lack of attention on ensemble-based attacks for dense prediction models such as object detection[26, 27]. Meanwhile, most of the existing ensemble model works[35, 36, 37] simply fuse the outputs of all models, neglecting the differences in model structures, which leads to adversarial optimization mostly heading in a single direction, making it difficult for adversarial examples to converge into a form that is effective for all target models. Recently, Cai et al.[17] proposed an attack algorithm based on transfer and query combination, demonstrating that adjusting the weights of the ensemble according to the victim model can further improve the attack performance. The universal adversarial perturbations they obtained can successfully attack various object detection models, such as YOLOv3 and Faster R-CNN, but can only be applied to digital images. Our research focused on the application of ensemble models in adversarial patches, adjusting weight coefficients based on energy loss feedback from different victim models to generate effective patches for all models.
@@ III-A Problem Definitions
In contrast to traditional methods of generating adversarial samples, adversarial patches are like stickers that are placed over specific local regions of an image. In order to investigate the generation mechanism of adversarial patches, we introduce the concept of the energy model and redefine the process of generating adversarial patches for object detection models from an energy perspective. In neural networks, "energy" is an abstract concept that does not directly correspond to energy in physics. In an energy model, "energy" refers to the value of the energy function, which is a description of the system state. When designing adversarial examples, an energy function can be used to quantify the model’s response or output distribution. By finding input perturbations that reduce the energy of model outputs (or increase them towards non-target classes), effective adversarial examples can be generated. Here, "energy" can be understood as a measure of the model’s confidence or certainty in specific outputs. By optimizing the energy function, attackers can devise inputs that lead to high uncertainty or errors in the model.
The energy function is usually written as $E(X,Y)$, which is used to measure the compatibility of ${X}$ and ${Y}$. It can also be understood as whether ${X,Y}$ matches or not, and the smaller the energy, the higher the match. The construction and training of an energy-based model requires the following parts: firstly, the structure of the energy function is determined, which we define as $\psi=E(G,X,Y),G\in\mathcal{G}$. When ${X,Y}$ are real vectors, ${\mathbf{\psi}}$ can be a simple linear combination, or a set of neural network architectures and weights. When ${X,Y}$ are variable images, sequences of symbols, vectors, or more complex structured objects, ${\mathbf{\varepsilon}}$ can represent a more complex class of functions. Thus, given data $S=\{(X^{i},Y^{i}):i=1...N\}$, we construct the energy function $L(E,S)$ to evaluate how good the current energy output is. For simplicity, this is usually denoted as $L(G,S)$. The purpose of the energy function is to find the parameter ${G}$ that minimizes the energy loss, i.e., $G^{{}^{*}}=\min_{G\in\mathcal{G}}L(G,S)$.
Given a clean sample ${X}$ with label ${Y}$, we aim to optimize the adversarial patch such that the person in the image evades recognition by the target detector. In contrast to traditional methods for generating adversarial samples, adversarial patches are like stickers placed on specific localized regions of an image. Therefore, we add adversarial patches to the target objects in the input image by means of a transformation function ${T(\cdot)}$. In order to investigate the generation mechanism of adversarial patches, we introduce the concept of energy model and redefine the process of generating adversarial patches for target detection models from an energy perspective. First, adversarial training is different from the training process of target detection models or energy-based models. The size and distribution of pixel values in the adversarial patch can be viewed as the parameters ${G}$ that need to be updated. Our goal is to construct a suitable energy function to find the optimal adversarial patch minimizing the energy, i.e., $p^{*}=\min_{p}L_{\mathrm{energy}}(p,S)$. The energy function for constructing adversarial patch on energy-based model is defined by the following equation:
@@
The training set ${S}$ is comprised of tuples ${\{({X^{i}},{Y^{i}});i=1...,N\}}$, which ${{X^{i}}}$ represents the input training example and ${{Y^{i}}}$ denotes its corresponding ground-truth label. ${p\in{\mathbb{R}^{H\times W\times 3}}}$ denotes the adversarial patch, and ${R(p)={\left\|p\right\|_{\mathbb{P}}}{\text{ = (}}\sum\limits_{i,j}{{{\left|{{p_{(i,j)}}}\right|}^{\mathbb{P}}}}{{\text{)}}^{1/\mathbb{P}}}}$ is the regularization term that restricts the size of each pixel ${{p_{(i,j)}}}$ within the adversarial patch by imposing a ${{l_{\mathbb{P}}}}$-norm constraint. During the adversarial patch training process, we optimize the confidence scores of candidate bounding boxes belonging to the “person” category, which are used as the objective for the optimization algorithm (i.e., the output of ${f(Y,T(p,X))}$). The energy function incorporates the confidence scores of all candidate boxes as part of the optimization process. The adversarial patch serves as the parameter to be updated iteratively until the energy of the target category is minimized through the optimization process.
@@ III-B Generating Adversarial Patch
Transformation Function: To facilitate the generalization of the adversarial patch from the digital to the physical world, we need to consider the interference of physical world factors on the patches during the training process. Therefore, based on the idea of Expectation over Transformation(EoT)[39] framework, we design the function ${{T}(\cdot)}$ to achieve random rotation and translation of the adversarial patch within a certain range to simulate the transformation of the shooting perspective in the physical world. At the same time, we propose the function ${{T}_{l}(\cdot)}$ to simulate the lighting transformation in the real world by adding random noise to the whole image and adjusting the brightness and contrast.In real-world physical environments, the scale of the patch tends to decrease as distance increases and the resulting imagery becomes more blurred. We introduce a novel distance transformation function. We employ a Gaussian kernel to blur the image, ensuring that the smaller the target object becomes, the larger the convolution kernel becomes, resulting in a more blurred adversarial patch image. Our goal is to develop an adversarial patch pattern that can be printed on a T-shirt to enable a person to evade object detection models. To achieve this goal, we introduce the Thin Plate Splines(TPS)[40] approach to simulate non-rigid deformations such as folds on the T-shirt’s surface. Using the aforementioned transformation methods, we devise a transformation function ${{T_{p}}(\cdot)}$ to apply the adversarial patch to the target image,
$\begin{gathered}{X_{adv}}={T_{p}}(p,X)=\hfill\\ {T_{l}}(((1-{M_{p}})\cdot X+{T}({T_{d}}({M_{p}}\cdot{T_{TPS}}(p)))))\hfill\\ \end{gathered}$ (3)
${X_{adv}}$ is the image after applying the counter patch, ${{M_{p}}\in{[0,1]^{d}}}$ represents the mask region of the adversarial patch on the target image, which is determined based on the bounding box obtained from the label. The ${{T_{p}}(\cdot)}$ function encompasses four main types of transformations, of which the first one involves applying TPS transformation ${T_{TPS}(\cdot)}$ to the generated adversarial patch. This serves to introduce distortions, thereby simulating non-rigid deformations on clothing surfaces. Subsequently, based on the size of the bounding box in the input image, the distance is estimated and a distance transformation function ${{T_{d}}(\cdot)}$ is applied to the generated adversarial patch. This enables the patch image to be blurred to varying degrees.The perspective transformation function ${{T}(\cdot)}$ is then used to simulate the perspective transformation during shooting in the physical world. In the last step, the illumination transformation ${T_{l}(\cdot)}$ is applied to the whole image to simulate the illumination transformation phenomenon in the physical world to further enhance the robustness of the adversarial patch.
@@
Adversarial Loss: For contemporary object detection models, it is crucial to design an appropriate energy function that can effectively deactivate the detection of the person category by minimizing the energy. Given images ${X}$ and the ground-truth labels ${Y}$, contemporary object detectors take this information as input, and utilize the detector’s prediction function ${{f}(\cdot)}$ to generate bounding boxes’ information. The information includes the information of the position of the candidate box, the confidence scores of the target, and the scores of each category. The candidates with confidence scores exceeding a predefined threshold are selected as the predicted results. Consequently, we incorporate the expected confidence scores of the output as a part of the energy function for adversarial patch optimization, aiming to reduce the target’s confidence scores by optimizing the adversarial patch. Specifically, we begin by randomly initializing the adversarial patch and utilizing the transformation function ${{T_{p}}(\cdot)}$ to apply the patch to the original image, which serves as the input for the object detection models. With the function ${{f}(\cdot)}$, we can get the information about the ${N}$ candidate boxes output by the detector, and ${f_{j}}$ denotes the information of the ${j}$ candidate box from output. Previous research incorporated the maximum confidence score of the target category, which was output by the detection network, as a constituent element in the energy function. Unlike previous work[18, 41], we consider all ${N}$ bounding boxes output by the detection model as optimization targets during each iteration, taking into account changes of all confidence scores comprehensively. To reduce the computational cost, we select candidate boxes whose confidence scores for the category “person” are higher than the set threshold value ${\mu}$ as attack objects. Subsequently, we continually optimize and update the patch pixels through gradient descent. In the case of one-stage object detection models, such as YOLOv2, YOLOv3, the detection network directly outputs all bounding boxes and associated category information. We integrate confidence scores of candidate boxes relevant to the “person” category as a constituent of the energy function. For two-stage object detection models like Faster R-CNN, person-related bounding boxes are extracted through Region Proposal Network(RPN), incorporating both positive and negative sample information. Positive candidate boxes with confidence scores surpassing a threshold are selectively chosen as attack targets. Hence, we compute the confidence score of all bounding boxes in the adversarial sample that exceeds the threshold ${\mu}$ of the target detection network. This score is merged into the energy function used for training the adversarial patch. Through iterative optimization, the patch is fine-tuned to minimize the total energy loss, as defined in Eq.4. The generative algorithm, designed based on the energy function, optimizes the adversarial patch during training, thereby reducing energy and causing a malfunction in the target detector’s person detection capability. By training a highly complex adversarial patch, the energy of the remaining categories is elevated, leading to a reduction in the energy associated with the person category. Sec.IV-E of the experiment provides a more detailed explanation regarding the impact of the adversarial patch on the energy levels of various categories.
@@ III-C Dynamically Optimized Ensemble Model
To achieve person invisibility across multiple object detection models, the generated adversarial patch effectively impacts the predictive information of all detection algorithms, resulting in a significant reduction of the energy associated with the person category in the output. For ${K}$ object detection models, ${f(\cdot)}$ denotes its outputs and ${\mathbf{w}=[{{\text{w}}_{1}},{{\text{w}}_{2}},...,{{\text{w}}_{K}}]}$ denotes the weights used to adjust different models, subject to constraint ${\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}}=1}$. In traditional approaches, the training is usually done by manually adjusting each value, or by directly averaging the values. The ensemble model for the adversarial patch can be expressed as follows,
$\min_{p}\sum_{i=1}^{K}\mathrm{w}_{i}\cdot\left(\mathbb{E}(\sum_{j=1}^{N}\max(f_{j}^{i}(Y,T_{p}(p,X)),\mu)\right)$ (5)
${f_{j}^{i}(\cdot)}$ denotes the information of the ${j\mathrm{-th}}$ candidate box output in the ${i\mathrm{-th}}$ detection model.The attack is achieved by optimizing the adversarial patches so that the outputs of all models are reduced. However, it is challenging to achieve an ideal balance point for the weights in object detection models with vastly different structures such as YOLO and Faster R-CNN, and overfitting is therefore prone to occur during the training process. Sec.IV-G of this paper provides a detailed analysis of this phenomenon. To address the aforementioned challenges, we propose a dynamically optimized ensemble model for multiple object detection models based on the Min-Max adversarial training approach. During the training process, the proportions of each detection model’s contribution to the overall energy output are dynamically adjusted through parameter tuning to achieve the universality of the adversarial patch, as calculated in Eq.6. The training procedure involves an outer minimization phase, with fixed parameters ${\mathbf{w}}$, where the adversarial patch is optimized through backpropagation to minimize the energy output of the ensemble model. And an inner maximization phase is performed with a fixed adversarial patch, using projected gradient descent to balance the weights ${\mathbf{w}}$ of different models and achieve the maximum energy output of the ensemble model. However, excessive parameter fluctuations can lead to oscillations during the training process, making it difficult to converge to an optimal adversarial patch. On the other hand, if the parameter fluctuations are too small, the loss may stop decreasing after reaching a certain point, failing to reach the desired level. In order to stabilize the process of the weights ${\mathbf{w}}$, the dynamic rate adjustment parameter $\nu$ is increased to control the speed of ${\mathbf{w}}$ change. The value of the parameter $\nu$ was verified in Sec.IV-A.
@@ IV-A Implementation Details
To conduct experiments on digital image attacks, we utilized the Inria Person dataset[44] consisting of 614 person images for training and 288 images for testing purposes. This dataset had been used extensively in research on adversarial attacks against the person category, such as the work of Thys et al.[18] and Hu et al.[21]. To evaluate the attack effectiveness and transferability of adversarial patches, we first measure their impact by computing the Average Precision (AP) values of the images with and without the patches. The larger the drop in AP, the stronger the attack effectiveness. In the physical experiments, to better evaluate the attack effectiveness of adversarial patches, we introduced the evaluation metric Attack Success Rate (ASR)[20]. ASR is defined as the proportion of images successfully attacked among all test images. Considering the existing baseline work, most of them are based on YOLOv2, Faster R-CNN for verification. Therefore, for a fair comparison, we designed our adversarial patches to attack mainly the mainstream single-stage object detection models, such as YOLOv2, YOLOv3, and YOLOv4. For more complex two-stage object detection models, we chose the Faster R-CNN detection model based on the backbone of the VGG16 and ResNet50 architectures. All detection models were trained on the COCO dataset, with the output confidence threshold set to 0.5 and the Non-Maximum Suppression (NMS) threshold to 0.4 during the inference stage. Due to laboratory arithmetic limitations, we only validated 2-3 combinations of target detectors in the ensemble model. Therefor, we consider three different sets of object detection models to design ensemble models for the adversarial patches:
@@
The entire algorithm is designed in PyTorch and trained on an NVIDIA RTX 3090 24GB GPU cluster. The adversarial patches are set to a size of $3\times 300\times 300$, optimized using the Adam[45] optimizer, and the learning rate for external minimization optimization is set to 0.03. We determined that the range of $\alpha$ parameter is in [0,0.05] and the range of $\beta$ parameter is in [0,0.5] based on the experience accumulated in previous related work. Then, by analyzing the generated adversarial images, we found that the gap between the patch when it is brighter and the printed effect is bigger, and the $\alpha$ parameter needs to be tuned down. If the noise on the patch is larger, it indicates that the image is not smooth enough and the parameter $\beta$ needs to be increased. Based on the above phenomenon, we determine that the parameters $\alpha=0.01$, $\beta=0.165$, $\mu=0.4$ and the regularization coefficient was set $\gamma=0.1$. The internal dynamic rate adjustment parameter $\nu$ was determined based on experimental results on DOEPatch(YF), as shown in Fig.4. Validation showed that with $\nu=0.78$, the adversarial patch achieved good attack effects on both YOLOv2 and YOLOv3.